Consider an economy described by the following equations:

C=800+0,8Y

I=400-2000i

G=120

NX=60-50-0,05Y

MS=2000

MD=0,2Y-2000i

Assuming that the interest rate equals i*=0,1, what is the level of equilibrium output?


1
Expert's answer
2021-06-08T12:17:05-0400

equilibrium output (Y): AD(Aggregate demand) = AS(Aggregate supply)

AD = C + I + G + NX

i = 0.1 therefore: I (investment) = 400 -2000(0.1)

I = 200

AD = 800+0.8Y+200+120+60-50-0.05Y

AD = 1130 + 0.75Y

AS = Y*


At money market equilibrium: MD = MS

0.2Y - 2000i = 2000

i = 0.1

0.2Y - 2000(0.1) = 2000

0.2Y - 200 = 2000

0.2Y = 2200

Y* = 11000

At Equilibrium output: AD = AS

Therefore: 1130 + 0.75Y = 11000

0.75Y = 9870

Y = 13160

Equilibrium output = 13160